OpenStudy (yonatanbesrat):

matty jogs 9 km/hr compute matty's speed in m/s

OpenStudy (okdutchman7):

Convert km/hr to m/s

OpenStudy (okdutchman7):

\[\frac{ 9km }{ 1hr }\times \frac{ 1000m }{ 1km }\times \frac{ 1hr }{ 60\min }\times \frac{ 1\min }{ 60\sec }\]

OpenStudy (okdutchman7):

All you have to do is multiply those together. All the units cancel out. That's why it's called a unit multiplier.